OpenStudy (anonymous):

1. How can we solve linear equations in one variable? 2. What are the similarities and differences in solving equations and inequalities? 3. What are the steps in solving an absolute value equation? 4. How is solving a literal equation different from solving a one- or two-step equation?

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

1. How can we solve linear equations in one variable? Um, I don't think this question makes any sense..can you check again?

OpenStudy (anonymous):

yea thats what it say my teacher made it

OpenStudy (anonymous):

If it means linear equations with one variable then we just isolate the variable on one side of the equation.

OpenStudy (anonymous):

oh ok thats the first one .

OpenStudy (anonymous):

2. What are the similarities and differences in solving equations and inequalities? They're basically the same..except in equations there are equal signs, and in inequalities there are inequality signs such as <, >, \(\le\), and \(\ge). And inequalities have a few more rules.

OpenStudy (anonymous):

and \(\ge\)*

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3. What are the steps in solving an absolute value equation? In absolute value equations, we create two equations. One will be the original without the absolute value signs, and one will be multiplying -1 into the absolute value bracket. Then we solve both.

OpenStudy (anonymous):

Okat Thanks :)

OpenStudy (anonymous):

"4. How is solving a literal equation different from solving a one- or two-step equation?" In Literal equations, we have to first convert it to a one-step or two-step equation.

OpenStudy (anonymous):

I believe that's all..

OpenStudy (anonymous):

okay thanks hun . ill let u know if i need help with anything else